OpenStudy (anonymous):

Solve the following equation. 4+2y=9+2y A.4 B.-4 C.infinitely many solutions D. no solution WILL GIVE MEDAL FOR CORRECT ANSWER!!!

OpenStudy (anonymous):

D

OpenStudy (anonymous):

2y-2y=9-4 0=5 It is not possible so no sollution D

OpenStudy (anonymous):

4+2y=9+2y 4=9 the answer is D. as there cant be a solution.

OpenStudy (anonymous):

WHen you get your answer it is 4=9. that is a no solution